OpenStudy (anonymous):

The sum of two numbers is 86. One of the numbers is five less than three times the other. What are the two numbers?

OpenStudy (anonymous):

x+y=86 x is five less than three times the other (x+5=3y)

OpenStudy (anonymous):

easiest way to solve this would be ot move the 5 to other side and sub in

OpenStudy (anonymous):

X+y=86 x-5=3y x=5+3y 5+3y+y=86 4y=86-5 4y=81 Y=81/4 Y=20.25 X+y=86 X+20.25=86 X=86-20.25

OpenStudy (anonymous):

so i seem to not understand why you make it x+5=3y? i was reading it asx-5=3y

OpenStudy (anonymous):

so it says that one of the numbers is 5 less than 3 times the other number. So that means x= 3y-5. This shows that x is 5 less than whatever 3y is, you get the 3y first, then subtract 5.

OpenStudy (anonymous):

ok thanks that is a big help

OpenStudy (anonymous):

Hope that helps make sense!